OpenStudy (anonymous):

5x-3y+2z=3 2x+4y-z=7 x-11y+4z=3

jhonyy9 (jhonyy9):

5x-3y+2z=3 2x+4y-z=7 multiply by 2 and assum with first equation x-11y+4z=3 5x-3y+2z=3 4x+8y-2z=14 -------------- 9x+5y 0 =17 2x+4y-z=7 multiply by 4 and assum with 3rd equation 8x+14y-4z=28 x-11y+4z=3 --------------- 9x+3y 0 =31 9x+5y=17 9x+3y=31 multiply by -1 and assum 9x+5y=17 -9x-3y=-31 ------------- 0 2y= -14 so y= -14/2 = -7 y=-7 9x+5(-7)=17 9x-35=17 9x=52 x=52/9 2x+4y-z=7 2(52/9)+4(-7)-z=7 104/9 -28-z=7 104 -252 -9z =63 -148 -9z =63 -148 -63=9z -211=9z z= -211/9 x= 52/9 y= -7 z= -211/9